OpenStudy (anonymous):

The radian measure of an angle that is 105(degrees) 32' 45'' is _____.

OpenStudy (anonymous):

\[\left(105 +\frac{32}{60}+\frac{45}{3600}\right)\left(\frac{\pi }{180}\right)=\frac{75993}{4 \pi }=1.84212 \]